Short Report – QA/QC Meeting July 2018

Meeting Date: 2018-07-12

Participants / Excused

LH, EPL, PM, IF, RM, KF, WE, MS, IB (9) / SB, MG, TB

Talking Points


  1. If you are (Tech-)SRP, please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
  2. SRP & Tech-SRP: please prepare short statement about your site and send to to LH (max. 2 sentences)
  3. Currently, automatic fluxes are calculated each day for the last 14 days. Most recent flux results are shown here, more information about the automatic calculation schedule (i.e. what is when calculated) is given here (deprecated). Auto-fluxes worked well for the last days, currently for sites CH-AWS, CH-CHA, CH-DAV (ETH EC files are used, not ICOS EC files), CH-FRU, CH-LAE and CH-OE2.
  4. Online QA/QC checks started for CH-DAV, CH-FRU, CH-CHA, CH-LAE and CH-OE2: Link Next step is to implement this check for CH-INO(?), CH-LAS(?) and CH-AWS, please contact LH.
  5. You can check the most recent eddy covariance raw data plots in the following folder on our group drive (P), updated every 2 days for your site:


  • Regarding GIN: RSS feed needed so that all news from GIN are automatically displayed on Swiss FluxNet homepage. DO will contact LH or LH will contact DO.
  • Reminder: the new Graphical Meteo Screening Tool by JE:
  • Roadmap

    1. [IN PROGRESS] SRP QA/QC Weekly checks in online checklists (instead of the server Excel sheet) are now implemented for CH-DAV, CH-FRU, CH-CHA, CH-LAE and CH-OE2.
    2. [IN PROGRESS] Subsrcibe to your site RSS feed to get relevant news automatically. Please contact LH for subscribing to the feed so you can get the newest Swiss FluxNet news for your site. Active subscriptions at the moment: SRPs: CH-CHA (IF), CH-OE2 (RM), CH-DAV (MG), CH-FRU (LH); T-SRPs: CH-DAV (PM), CH-CHA (PL), CH-FRU (PL).
    3. [IN PROGRESS] On-demand flux calculations on grassland server dedicated ETH server instead of RDS. In progress.
    4. [UPCOMING] CH-INO will be added to these reports. LH and RM set up the flux processing soon.
    5. [UPCOMING] QV File Reports: checks available files on server. As soon as ICOS allows.
    6. [UPCOMING] T-SRP QA/QC Weekly checks in online checklists. T-SRPs will contact LH for checklist items.
    7. [DONE] Automatic fluxes will be activated again (goal is April 2018), priority conflicts w/ ICOS. All sites now active.
    8. [DONE] Activate site-specific RSS feeds. RSS feeds are now active.

    Site Reports


    SRP (LH, fluxes): High CO2 uptake over the last month, EC system ran well without problems.
    Tech-SRP (IB): The CO2/H2O analyzer, the 3D ultrasonic anemometer and the weather station are running normally. The adaptive mast and the snowheight sensor are currently out of order, but not needed during summer time.


    SRP (IF): The cable from the analyzer box of the Irga was not plugged in properly, like that there was only an intermittent connection to the logger, resulting in a data loss of 3 consecutive days (23-26 June) and intermittent data gaps until July 4 when it got fixed. Slurry attack on the radiation and other meteo sensors on June 26. The QCL will stay at Aerodyne not longer than until November but might come back earlier as soon as they solve the current drop issue.
    Tech-SRP (PL): –


    SRP (MG, LH): Except for the data gap in June (memory problem?, see PM below) the EC system runs fine and fluxes look good. We implemented a new script that empties the OK folder on the LANNER automatically every month (?) to make enough room for new incoming data. Currently fluxes are only calculated from the ETH EC files, but fluxes from the ICOS EC files will also be implemented later.
    Tech-SRP (PM): Data acquisition computer ran out of memory. This is probably the reason for a EC data gap from 19.06.2018-26.06.2018. Lots of work has been done at the ICOS valve switching system.


    SRP (LH): No problems found, EC and meteo data are complete and look good.
    Tech-SRP (PL): –


    SRP (EPL): Both CH-LAE and CH-LAS have been running well within the last month. We still need to find a way to remove or flag the fluxes calculated when there was no flow for the LI-7200 at the beginning of June.
    Tech-SRP (TB): –


    SRP (RM): Everthing is running and looking good. IRGA needs a calibration in August.
    Tech-SRP (IB): The CO2 analyzer and 3D ultrasonic anemometer are running normally. Also the logger.


    SRP (RM): checks will be implemented in the coming weeks
    T-SRP (IB):
    Laser spectrometer, CO2 analyzer and 3D ultrasonic anemometer are running normally. The station is currently placed in Oensingen.


    You are here: Short Report - QA/QC Meeting July 2018


    Last Updated on 14 Feb 2024 12:20

    Leave a Reply

    You must be logged in to post a comment.